OpenStudy (supersparky277):

Find the fifth term in the arithmetic sequence that begins at 6.7 and has a common difference of -20.4.

OpenStudy (sshayer):

a=6.7 d=-20.4 n=5 \[t _{n }=a+(n-1)d \]

OpenStudy (calculusxy):

\(a_n = a_1 + (n-1)d\) where \(a_n\) = the term \(a_1\) = the first term \(n\) = the position of the term ( first, second, etc.) \(d\) = common difference in this case \(a_1\) = 6.7 \(d\) = -20.4 \(a_5 = 6.7 + ((5 - 1) \times-20.4)\) Evaluate that to find what \(a_5\) (or the fifth term) is.